OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

Directrix (directrix):

y = x + 6 y = 2x Solve these by substitution. x + 6 = 2x -x = - 6 x = 6 y = 2x y = 2(6) = 12 Solution ( 6, 12 )